# Improved Margin Generalization Bounds for Voting Classifiers

### v1

@@ 1 Introduction
Creating ensembles of classifiers is a classic technique in machine learning, and thus studying generalization of such ensembles is natural. An example of the success of ensembles of classifiers is boosting algorithms, which are one of the pillars of classic machine learning, often significantly improving the accuracy of a base learning algorithm by creating an ensemble of multiple base classifiers/hypotheses. More formally, consider binary classification over an input domain ${\mathcal{X}}$ and let ${\mathcal{A}}$ be a base learning algorithm that on a (possibly weighted) training sequence $S\in({\mathcal{X}}\times\{-1,1\})^{*}$, returns a hypothesis $h_{S}\in\{-1,1\}^{{\mathcal{X}}}$ better than guessing on the weighted training sequence. Boosting algorithms, like AdaBoost [Freund and Schapire, 1997] then iteratively invokes ${\mathcal{A}}$ on reweighed version of the training sequence $S$ to produce hypotheses $h_{1},\dots,h_{t}$. In each step, the training sequence $S$ is weighed to put more emphasis on data points that $h_{1},\dots,h_{t-1}$ misclassified, forcing $h_{t}$ to focus on these points. The obtained hypotheses are finally combined via a weighted majority vote to obtain the voting classifier $f(x)=\operatorname{\mathrm{sign}}(\sum_{i}\alpha_{i}h_{i}(x))$ with $\alpha_{i}>0$ for all $i$.
Much research has gone into understanding the impressive performance of boosting algorithms and voting classifiers, with one particularly influential line of work focusing on so-called margins [Schapire et al., 1998]. Given a voting classifiers $f(x)=\operatorname{\mathrm{sign}}(\sum_{i}\alpha_{i}h_{i}(x))$, the margin of $f$ on a data point $(x,y)\in{\mathcal{X}}\times\{-1,1\}$ is then defined as
@@
Observe that the margin is a number in $[-1,1]$. The margin is $1$ if all hypotheses $h_{i}$ in $f$ agree and are correct, it is $-1$ if they agree and are incorrect, and it is 00 if there is a 50-50 weighted split in their predictions for the label of $x$. The margins can thus be thought of as a (signed) certainty in the prediction made by $f$.
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f):=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[f({\mathbf{x}})\neq{\mathbf{y}}]=\operatorname{\mathcal{L}}^{\gamma}_{\mathbf{S}}(f)+O\left(\sqrt{\frac{\ln(|{\mathcal{H}}|)\ln m}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}}\right).$ (1)
@@
This matches the upper bound in Eq. 3 up to the gap between $\sqrt{\ln(1/\tau)}\approx\sqrt{\ln(1/\operatorname{\mathcal{L}}^{\gamma}_{{\mathbf{S}}}(f))}$ and $\ln m$, improving by a $\sqrt{\ln(1/\tau)}$ factor over a previous lower bound by [Grønlund et al., 2019]. Note that we have simplified the lower bound slightly, as the true statement would have $\ln m$ replaced by $\ln(\gamma^{2}m/\ln N)$. A similar substitution of $\ln m$ by $\ln(\gamma^{2}m/\ln(|{\mathcal{H}}|))$ in the upper bound Eq. 3 is also possible.
@@
One of the prime motivations for studying generalization bounds for large margin voting classifiers, is their application to boosting algorithms. When studying boosting theoretically, we typically use the framework of weak to strong learning by [Kearns and Valiant, 1988, Kearns and Valiant, 1994]. Let $t\in\{-1,1\}^{\mathcal{X}}$ be an unknown target concept assigning labels $t(x)$ to samples $x\in{\mathcal{X}}$. For a distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, let ${\mathcal{D}}_{t}$ be the distribution over ${\mathcal{X}}\times\{-1,1\}$ obtained by drawing a sample ${\mathbf{x}}\sim{\mathcal{D}}$ and returning the pair $({\mathbf{x}},t({\mathbf{x}}))$.
A $\gamma$-weak learner ${\mathcal{W}}$, is a learning algorithm that for any distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, when given $m_{0}$ i.i.d. samples $({\mathbf{x}}_{i},t({\mathbf{x}}_{i}))\sim{\mathcal{D}}_{t}$, ${\mathcal{W}}$ produces with probability at least $1-\delta_{0}$ a hypothesis $h$ with $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(h)=\operatorname*{\mathbb{P}}_{({\mathbf{x}},t({\mathbf{x}}))\sim{\mathcal{D}}_{t}}[h({\mathbf{x}})\neq t({\mathbf{x}})]\leq 1/2-\gamma$. Here $m_{0}$ and $\delta_{0}$ are constants. A strong learner in contrast, is a learning algorithm that for any distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, when given $m(\varepsilon,\delta)$ i.i.d. samples from ${\mathcal{D}}_{t}$, it produces with probability at least $1-\delta$ a hypothesis with $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(h)\leq\varepsilon$. A strong learner thus obtains arbitrarily high accuracy when given enough training data.
AdaBoost [Freund and Schapire, 1997] is the most famous algorithm for constructing a weak learner from a strong learner. Concretely, it can be shown that if AdaBoost is run for $O(\gamma^{-2}\ln m)$ iterations, then it produces a voting classifier $f$ with margin $\Omega(\gamma)$ on all samples in a given training sequence $S$ with $|S|=m$ [Schapire and Freund, 2012] [Theorem 5.8]. If the weak learner/base learning algorithm always returns hypotheses from a hypothesis class ${\mathcal{H}}$ of VC-dimension $d$, this allows us to use our new generalization bound in Theorem 1 to conclude
@@
For any $\gamma$-weak learner ${\mathcal{W}}$ using a hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ of VC-dimension $d$, distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $t$, failure parameter $10<\delta<$ 1 and any constant $10<\varepsilon<$ 1, it holds with probability at least $1-\delta$ over ${\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}$, that the voting classifier $f$ produced by AdaBoost on ${\mathbf{S}}$ with weak learner ${\mathcal{W}}$ has
$\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(f)=O\left(\frac{d\operatorname{\Gamma}(\gamma^{2}m/d)}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}\right),$
@@
The previous best known generalization bound for AdaBoost followed from the margin generalization bound Eq. 5 and was $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(f)=O\left(\frac{d\ln(m/d)\ln m}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}\right)$. Moreover, our new bound is tight up to a $\ln^{2}(\ln(\gamma^{2}m/d))$ factor as demonstrated by a lower bound of [Høgsgaard et al., 2023] showing that for $c^{-1}\sqrt{d/m}<\gamma<c$ for sufficiently small constant $c>0$ and VC-dimension $d=\Omega(\ln(1/\gamma))$, there is a data distribution ${\mathcal{D}}$, a weak learner ${\mathcal{W}}$ using a hypothesis class of VC-dimension $d$, and a concept $t$, such that AdaBoost run with ${\mathcal{W}}$ has $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(f)=\Omega\left(\frac{d\ln(\gamma^{2}m/d)}{\gamma^{2}m}\right)$, with constant probability over a training sequence ${\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}$.
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(f)=\Theta\left(\frac{d}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}\right).$ (6)
In light of the lower bound above for AdaBoost, this implies that AdaBoost is not an optimal weak to strong learner. However, several optimal weak to strong learning algorithms have been developed. In [Larsen and Ritzert, 2022], the authors gave the first such algorithm. This algorithm uses the sub-sampling idea of [Hanneke, 2016] from optimal realizable PAC learning and runs AdaBoost on $m^{\lg_{4}3}$ many sub-samples $S_{i}\subset S$ of the training data. It combines the produced voting classifiers by taking a majority vote among their predictions, i.e. a majority-of-majorities. Since each $S_{i}$ has $|S_{i}|=\Omega(m)$, this slows down AdaBoost by a factor $m^{\lg_{4}3}\approx m^{0.79}$. Later work by [Larsen, 2023] showed that running Bagging [Breiman, 1996] to draw $O(\ln(m/\delta))$ many random sub-samples $S_{i}$ from $S$ and running AdaBoost on each also results in an optimal generalization error matching Eq. 6, thus reducing the computational overhead to a logarithmic factor. Finally, a recent work by [Høgsgaard et al., 2024] built on a Majority-of-3 result in realizable PAC learning [Aden-Ali et al., 2024] to show that simply partitioning a training sequence into $5$ disjoint pieces of $m/5$ samples each, and outputting a majority vote among voting classifiers trained with AdaBoost on each sub-sample, results in an optimal in-expectation error of $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(f)]=O(d/(\gamma^{2}m))$.
Our new generalization bound in Theorem 1 allows us to improve the Majority-of-5 algorithm to a more natural Majority-of-3
@@
[Follows from Theorem 11] For any $\gamma$-weak learner ${\mathcal{W}}$ using a hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ of VC-dimension $d$, distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, and concept $t$, it holds that the voting classifiers $f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}$ produced by AdaBoost on i.i.d. training sequences ${\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{t}^{m}$ with weak learner ${\mathcal{W}}$ satisfy
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{t}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))\right]=O\left(\frac{d}{\gamma^{2}m}\right).$
@@ 2.1 New Margin Generalization Bounds
where $\operatorname{\Gamma}(x)=\ln(x)\ln^{2}(\ln x)$.
Let us first introduce a more convenient way of representing voting classifiers. Recall that voting classifiers $f$ are of the form $f(x)=\operatorname{\mathrm{sign}}(\sum_{i}\alpha_{i}h_{i}(x))$ with all $\alpha_{i}>0$. Furthermore, the margin on a training example $(x,y)$ is defined as $y\sum_{i}\alpha_{i}h_{i}(x)/\sum_{j}\alpha_{j}$. To avoid the tedious normalization by $\sum_{j}\alpha_{j}$, we henceforth assume all voting classifiers have $\sum_{i}\alpha_{i}=1$. In addition, we will drop the $\operatorname{\mathrm{sign}}(\cdot)$ and instead write $f(x)=\sum_{i}\alpha_{i}h(x)$. In this way, we have that $\operatorname{\mathrm{sign}}(f(x))\neq y$ if and only if $yf(x)\leq 0$ (we define $\operatorname{\mathrm{sign}}(0)=0$). Furthermore, the margin is no more than $\gamma$ if and only if $yf(x)\leq\gamma$. We hence define $\operatorname{\Delta({\mathcal{H}})}$ as the set of all convex combinations $\sum_{i}\alpha_{i}h_{i}$ for $h_{i}\in{\mathcal{H}}$ (i.e. $\sum_{i}\alpha_{i}=1$ and $\alpha_{i}>0$) and refer to $\operatorname{\Delta({\mathcal{H}})}$ as the set of voting classifiers over ${\mathcal{H}}$. With this notation, we have $\operatorname{\mathcal{L}}^{\gamma}_{\mathcal{D}}(f)=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[{\mathbf{y}}f({\mathbf{x}})\leq\gamma]$, $\operatorname{\mathcal{L}}^{\gamma}_{S}(f)$ is the fraction of training examples $(x,y)\in S$ with $yf(x)\leq\gamma.$ For $\gamma=0$ we will use $\operatorname{\mathcal{L}}_{\mathcal{D}}(f)=\operatorname{\mathcal{L}}^{0}_{\mathcal{D}}(f)=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[{\mathbf{y}}f({\mathbf{x}})\leq\gamma]=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[\operatorname{\mathrm{sign}}(f({\mathbf{x}}))\not={\mathbf{y}}]$.
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)\leq\tau_{1}+O\left(\sqrt{\tau_{1}\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)}+\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)\right).$ (8)
@@
$\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}(f)\leq\tau_{1}+O\left(\sqrt{\tau_{1}\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)}+\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)\right).$ (9)
@@
To construct a $\gamma_{0}/2$ $\ell_{\infty}$-covering $N$ of ${\mathbf{X}}$ and union bound over it, we need the point set ${\mathbf{X}}$ to be fixed - however we still want to be able to show that an imbalance between $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma_{0}/2}(f^{\prime})$ and $\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}^{\gamma_{0}/2}(f^{\prime})$ for some $f^{\prime}\in N$ is highly unlikely. As in previous works, we employ the following way of viewing the sampling of ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$. First, we draw ${\mathbf{X}}\sim{\mathcal{D}}^{2m}$, consisting of $2m$ i.i.d. training examples from ${\mathcal{D}}$, and then let ${\mathbf{S}}$ be $m$ points drawn without replacement from ${\mathbf{X}},$ and ${\mathbf{S}}^{\prime}$ be the remaining $m$ points of ${\mathbf{X}},$ i.e., ${\mathbf{S}}^{\prime}={\mathbf{X}}\backslash{\mathbf{S}}^{\prime}.$ Taking this viewpoint of drawing ${\mathbf{S}}$ and ${\mathbf{S}}$ allows us to fix the realization $X$ of the points in ${\mathbf{X}},$ while still having which training examples ending up in ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ being random. This still allows us to argue that an imbalance between $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma_{0}/2}(f^{\prime})$ and $\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}^{\gamma_{0}/2}(f^{\prime})$ for some $f^{\prime}\in N$ is unlikely.

### v2

@@ 1 Introduction
Creating ensembles of classifiers is a classic technique in machine learning, and thus studying generalization of such ensembles is natural. An example of the success of ensembles of classifiers is boosting algorithms, which are one of the pillars of classic machine learning, often significantly improving the accuracy of a base learning algorithm by creating an ensemble of multiple base classifiers/hypotheses. More formally, consider binary classification over an input domain ${\mathcal{X}}$ and let ${\mathcal{A}}$ be a base learning algorithm that on a (possibly weighted) training sequence $S\in({\mathcal{X}}\times\{-1,1\})^{*}$, returns a hypothesis $h_{S}\in\{-1,1\}^{{\mathcal{X}}}$ better than guessing on the weighted training sequence. Boosting algorithms, like AdaBoost [Freund and Schapire, 1997] then iteratively invokes ${\mathcal{A}}$ on reweighed version of the training sequence $S$ to produce hypotheses $h_{1},\dots,h_{t}$. In each step, the training sequence $S$ is weighed to put more emphasis on data points that $h_{1},\dots,h_{t-1}$ misclassified, forcing $h_{t}$ to focus on these points. The obtained hypotheses are finally combined via a weighted majority vote to obtain the voting classifier $f(x)=\sum_{i}\alpha_{i}h_{i}(x)$ with $\alpha_{i}>0$ for all $i$. Much research has gone into understanding the impressive performance of boosting algorithms and voting classifiers, with one particularly influential line of work focusing on so-called margins [Schapire et al., 1998]. Given a voting classifiers $f(x)=\sum_{i}\alpha_{i}h_{i}(x)$, the margin of $f$ on a data point $(x,y)\in{\mathcal{X}}\times\{-1,1\}$ is then defined as
@@
Observe that the margin is a number in $[-1,1]$. The margin is $1$ if all hypotheses $h_{i}$ in $f$ agree and are correct, it is $-1$ if they agree and are incorrect, and it is 00 if there is a 50-50 weighted split in their predictions for the label of $x$. The margins can thus be thought of as a (signed) certainty in the prediction made by $f$. Naturally $\operatorname{\mathrm{sign}}(f(x))$ correctly predicts the label $y$ if and only if $\operatorname{margin}(f,(x,y))>0$, by taking $\operatorname{\mathrm{sign}}(0)=0.$
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f):=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[{\mathbf{y}}f({\mathbf{x}})\leq 0]=\operatorname{\mathcal{L}}^{\gamma}_{\mathbf{S}}(f)+O\left(\sqrt{\frac{\ln(|{\mathcal{H}}|)\ln m}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}}\right).$ (1)
@@
This matches the upper bound in Eq. 3 up to the gap between $\sqrt{\ln(1/\tau)}\approx\sqrt{\ln(1/\operatorname{\mathcal{L}}^{\gamma}_{{\mathbf{S}}}(f))}$ and $\ln m$, improving by a $\sqrt{\ln(1/\tau)}$ factor over a previous lower bound by [Grønlund et al., 2019]. Note that we have simplified the lower bound slightly, as the true statement would have $\ln m$ replaced by $\ln(\gamma^{2}m/\ln N)$. A similar substitution of $\ln m$ by $\ln(\gamma^{2}m/\ln(|{\mathcal{H}}|))$ in the upper bound Eq. 3 also seems possible with minor modifications to the previous proof.
@@
One of the prime motivations for studying generalization bounds for large margin voting classifiers, is their application to boosting algorithms. When studying boosting theoretically, we typically use the framework of weak to strong learning by [Kearns and Valiant, 1988, Kearns and Valiant, 1994]. Let $c^{*}\in\{-1,1\}^{\mathcal{X}}$ be an unknown target concept assigning labels $c^{*}(x)$ to samples $x\in{\mathcal{X}}$. For a distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, let ${\mathcal{D}}_{c^{*}}$ be the distribution over ${\mathcal{X}}\times\{-1,1\}$ obtained by drawing a sample ${\mathbf{x}}\sim{\mathcal{D}}$ and returning the pair $({\mathbf{x}},c^{*}({\mathbf{x}}))$.
A $\gamma$-weak learner ${\mathcal{W}}$, is a learning algorithm that for any distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, when given $m_{0}$ i.i.d. samples $({\mathbf{x}}_{i},c^{*}({\mathbf{x}}_{i}))\sim{\mathcal{D}}_{c^{*}}$, ${\mathcal{W}}$ produces with probability at least $1-\delta_{0}$ a hypothesis $h$ with $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(h)=\operatorname*{\mathbb{P}}_{({\mathbf{x}},c^{*}({\mathbf{x}}))\sim{\mathcal{D}}_{c^{*}}}[h({\mathbf{x}})\neq c^{*}({\mathbf{x}})]\leq 1/2-\gamma$. Here $m_{0}$ and $\delta_{0}$ are constants. A strong learner in contrast, is a learning algorithm that for any distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, when given $m(\varepsilon,\delta)$ i.i.d. samples from ${\mathcal{D}}_{c^{*}}$, it produces with probability at least $1-\delta$ a hypothesis with $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(h)\leq\varepsilon$. A strong learner thus obtains arbitrarily high accuracy when given enough training data. AdaBoost [Freund and Schapire, 1997] is the most famous algorithm for constructing a strong learner from a weak learner. Concretely, it can be shown that if AdaBoost is run for $O(\gamma^{-2}\ln m)$ iterations, then it produces a voting classifier $f$ with margin $\Omega(\gamma)$ on all samples in a given training sequence $S$ with $|S|=m$ [Schapire and Freund, 2012] [Theorem 5.8]. If the weak learner/base learning algorithm always returns hypotheses from a hypothesis class ${\mathcal{H}}$ of VC-dimension $d$, this allows us to use our new generalization bound in Theorem 1 to conclude
@@
For any $\gamma$-weak learner ${\mathcal{W}}$ using a hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ of VC-dimension $d$, distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, target concept $c^{*}$, failure parameter $10<\delta<$ 1 and any constant $10<\varepsilon<$ 1, it holds with probability at least $1-\delta$ over ${\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}$, that the voting classifier $f$ produced by AdaBoost on ${\mathbf{S}}$ with weak learner ${\mathcal{W}}$ has
$\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(f)=O\left(\frac{d\operatorname{\Gamma}(\gamma^{2}m/d)}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}\right),$
@@
The previous best known generalization bound for AdaBoost followed from the margin generalization bound Eq. 5 and was $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(f)=O\left(\frac{d\ln(m/d)\ln m}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}\right)$. Moreover, our new bound is tight up to a $\ln^{2}(\ln(\gamma^{2}m/d))$ factor as demonstrated by a lower bound of [Høgsgaard et al., 2023] showing that for $\gamma<c$ for sufficiently small constant $c>0$ and VC-dimension $d=\Omega(\ln(1/\gamma))$, and sample size $d\gamma^{-2}/16\leq m\leq\exp{\left(\exp{\left(d\right)}\right)},$ there is a data distribution ${\mathcal{D}}$, a weak learner ${\mathcal{W}}$ using a hypothesis class of VC-dimension $d$, and a concept $c^{*}$, such that AdaBoost run with ${\mathcal{W}}$ has $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(f)=\Omega\left(\frac{d\ln(\gamma^{2}m/d)}{\gamma^{2}m}\right)$, with constant probability over a training sequence ${\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}$.
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(f)=\Theta\left(\frac{d}{\gamma^{2}m}+\frac{\ln(1/\delta)}{m}\right).$ (6)
In light of the lower bound above for AdaBoost, this implies that AdaBoost is not an optimal weak to strong learner. However, several optimal weak to strong learning algorithms have been developed. In [Larsen and Ritzert, 2022], the authors gave the first such algorithm. This algorithm uses the sub-sampling idea of [Hanneke, 2016] from optimal realizable PAC learning and runs AdaBoost on $m^{\lg_{4}3}$ many sub-samples $S_{i}\subset S$ of the training data. It combines the produced voting classifiers by taking a majority vote among their predictions, i.e. a majority-of-majorities. Since each $S_{i}$ has $|S_{i}|=\Omega(m)$, this slows down AdaBoost by a factor $m^{\lg_{4}3}\approx m^{0.79}$. Later work by [Larsen, 2023] showed that running Bagging [Breiman, 1996] to draw $O(\ln(m/\delta))$ many random sub-samples $S_{i}$ from $S$ and running AdaBoost on each also results in an optimal generalization error matching Eq. 6, thus reducing the computational overhead to a logarithmic factor. Finally, a recent work by [Høgsgaard et al., 2024] built on a Majority-of-3 result in realizable PAC learning [Aden-Ali et al., 2024] to show that simply partitioning a training sequence into $5$ disjoint pieces of $m/5$ samples each, and outputting a majority vote among voting classifiers trained with AdaBoost on each sub-sample, results in an optimal in-expectation error of $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(f)]=O(d/(\gamma^{2}m))$.
Our new generalization bound in Theorem 1 allows us to improve the Majority-of-5 algorithm to a more natural Majority-of-3 of AdaBoost (we follow the AdaBoost implementation on page 5 [Schapire and Freund, 2012]). We here assume that Adaboost has access to a deterministic empirical $\gamma$-weak learner, where an empirical $\gamma$-weak learner, takes as input a sample ${\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{n}$ and a distribution ${\mathcal{D}}^{\prime}$ over ${\mathbf{S}}$ and returns a hypothesis $h$ such that $\operatorname{\mathcal{L}}_{{\mathcal{D}}^{\prime}}(h)\leq 1/2-\gamma.$ The existence of such a empirical $\gamma$-weak learner is given by the prior definition of a $\gamma$-weak learner.
@@
[Follows from Theorem 11] For any $\gamma$-weak learner ${\mathcal{W}}$ using a hypothesis class ${\mathcal{H}}\subseteq\{-1,1\}^{{\mathcal{X}}}$ of VC-dimension $d$, distribution ${\mathcal{D}}$ over ${\mathcal{X}}$, and concept $c^{*}$, it holds that the voting classifiers $f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}$ produced by AdaBoost on i.i.d. training sequences ${\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{c^{*}}^{m}$ with weak learner ${\mathcal{W}}$ satisfy
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{1},{\mathbf{S}}_{2},{\mathbf{S}}_{3}\sim{\mathcal{D}}_{c^{*}}^{m}}\left[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))\right]=O\left(\frac{d}{\gamma^{2}m}\right).$
@@ 2.1 New Margin Generalization Bounds
where $\operatorname{\Gamma}(x)=\ln(x)\ln^{2}(\ln x)$. Let us first introduce a more convenient way of representing voting classifiers. Recall that voting classifiers $f$ are of the form $f(x)=\sum_{i}\alpha_{i}h_{i}(x)$ with all $\alpha_{i}>0$. Furthermore, the margin on a training example $(x,y)$ is defined as $y\sum_{i}\alpha_{i}h_{i}(x)/\sum_{j}\alpha_{j}$. To avoid the tedious normalization by $\sum_{j}\alpha_{j}$, we henceforth assume all voting classifiers have $\sum_{i}\alpha_{i}=1$. In this way, we have that $\operatorname{\mathrm{sign}}(f(x))\neq y$ if and only if $yf(x)\leq 0$ (we define $\operatorname{\mathrm{sign}}(0)=0$). Furthermore, the margin is no more than $\gamma$ if and only if $yf(x)\leq\gamma$. We hence define $\operatorname{\Delta({\mathcal{H}})}$ as the set of all convex combinations $\sum_{i}\alpha_{i}h_{i}$ for $h_{i}\in{\mathcal{H}}$ (i.e. $\sum_{i}\alpha_{i}=1$ and $\alpha_{i}>0$) and refer to $\operatorname{\Delta({\mathcal{H}})}$ as the set of voting classifiers over ${\mathcal{H}}$. With this notation, we have $\operatorname{\mathcal{L}}^{\gamma}_{\mathcal{D}}(f)=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[{\mathbf{y}}f({\mathbf{x}})\leq\gamma]$, $\operatorname{\mathcal{L}}^{\gamma}_{S}(f)$ is the fraction of training examples $(x,y)\in S$ with $yf(x)\leq\gamma.$ For $\gamma=0$ we will use $\operatorname{\mathcal{L}}_{\mathcal{D}}(f)=\operatorname{\mathcal{L}}^{0}_{\mathcal{D}}(f)=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[{\mathbf{y}}f({\mathbf{x}})\leq 0]=\operatorname*{\mathbb{P}}_{({\mathbf{x}},{\mathbf{y}})\sim{\mathcal{D}}}[\operatorname{\mathrm{sign}}(f({\mathbf{x}}))\not={\mathbf{y}}]$.
@@
$\operatorname{\mathcal{L}}_{{\mathcal{D}}}(f)=\tau_{1}+O\left(\sqrt{\tau_{1}\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)}+\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)\right).$ (8)
@@
$\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}(f)=\tau_{1}+O\left(\sqrt{\tau_{1}\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)}+\left(\frac{d\operatorname{\Gamma}{\left(\frac{m\gamma_{0}^{2}}{d}\right)}}{\gamma_{0}^{2}m}+\frac{\ln{\left(\frac{1}{\delta}\right)}}{m}\right)\right).$ (9)
@@
To construct a $\gamma_{0}/2$ $\ell_{\infty}$-covering $N$ of ${\mathbf{X}}$ and union bound over it, we need the point set ${\mathbf{X}}$ to be fixed - however we still want to be able to show that an imbalance between $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma_{0}/2}(f^{\prime})$ and $\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}^{\gamma_{0}/2}(f^{\prime})$ for some $f^{\prime}\in N$ is highly unlikely. As in previous works, we employ the following way of viewing the sampling of ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$. First, we draw ${\mathbf{X}}\sim{\mathcal{D}}^{2m}$, consisting of $2m$ i.i.d. training examples from ${\mathcal{D}}$, and then let ${\mathbf{S}}$ be $m$ points drawn without replacement from ${\mathbf{X}},$ and ${\mathbf{S}}^{\prime}$ be the remaining $m$ points of ${\mathbf{X}},$ i.e., ${\mathbf{S}}^{\prime}={\mathbf{X}}\backslash{\mathbf{S}}^{\prime}.$ Taking this viewpoint of drawing ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ allows us to fix the realization $X$ of the points in ${\mathbf{X}},$ while still having which training examples ending up in ${\mathbf{S}}$ and ${\mathbf{S}}^{\prime}$ being random. This still allows us to argue that an imbalance between $\operatorname{\mathcal{L}}_{{\mathbf{S}}}^{\gamma_{0}/2}(f^{\prime})$ and $\operatorname{\mathcal{L}}_{{\mathbf{S}}^{\prime}}^{\gamma_{0}/2}(f^{\prime})$ for some $f^{\prime}\in N$ is unlikely.